Circular No. 8713

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006ci, 2006cj, 2006ck Three apparent supernovae have been found on unfiltered CCD frames:

2006ci by C. Colesanti, C. Jacques, E. Pimentel, and T. Napoleao (cf. IAUC 8658); 2006cj by R. Quimby, P. Mondol, and F. Castro (cf. IAUC 8622); and 2006ck by P. Luckas, O. Trondal, and M. Schwartz (cf. IAUC 8687).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ci May 18.20 18^{h}18^{m}28.^{s}67 −54^{o}41′45′′.6 15.1 24′′.7 W, 6′′.5 S
2006cj May 17.23 12 59 24.42 +28 20 51.7 18.1 4′′ E, 2′′ N
2006ck May 20.54 13 09 40.42 − 1 02 57.2 17.4 10′′.2 W, 3′′.0 S

Additional magnitudes for 2006ci in ESO 182-G10: Apr. 5.39 UT, [19.0; May 19.05, 15.2. Nothing is visible at the location of 2006ci on a red Palomar Survey image from 1995. Additional approximate magnitudes for 2006cj: 2004 Dec. 15 and 2006 Jan. 10, [19.3; May 18.24, 18.0; 19.28, 17.7 (poor conditions). Spectroscopy by S. Blondin et al. (cf. CBET 517) on May 22.32 shows that 2006cj is a type-Ia supernova just past maximum. Additional magnitudes for 2006ck in UGC 8238: 1989 Feb. 11, [21.0 (Digitized Sky Survey, red); Mar. 29.70, [18.0 (Luckas); May 22.17, 17.4. Blondin et al. report (cf. CBET 519) that a spectrogram taken on May 23.21 shows 2006ck to be a type-Ic supernova near maximum.

COMETS C/2006 D3–D6, C/2006 E3, C/2006 E4 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8712). C/2006 D3 and C/2006 D4 were described by K. Battams as tiny, stellar, and extremely faint. C/2006 D5 had a teardrop shape, reaching mag 5.2 on Feb. 26.738 UT at 11.8R_{⊙}. C/2006 D6 and C/2006 E3 appeared stellar, both reaching mag 4.6, on Mar. 1.358 at 11.7R_{⊙} and Mar. 13.399 at 11.4R_{⊙}, respectively. C/2006 E4 was extremely faint (mag ∼ 8.5) and diffuse.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 D3 Feb. 19.738 22^{h}27.^{m}8 −12^{o}07′ C3 SF 2006-J51
C/2006 D4  21.138 22 32.9 −11 31 C3 HS 2006-J51
C/2006 D5  26.279 22 58.3 − 9 59 C3 SH 2006-K09
C/2006 D6  28.691 23 11.0 − 9 16 C3 SF 2006-K09
C/2006 E3 Mar. 12.899 23 53.4 − 4 15 C3 HS 2006-K09
C/2006 E4  15.133 23 47.2 − 2 27 C2 TH 2006-K09
2006 May 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green